June 26, 2014 BCAL Commissioning

From GlueXWiki
Jump to: navigation, search

Video Conferencing Information

Meeting Time: 2:15 p.m. Eastern

  1. To join via a Web Browser, go to the page [1] https://bluejeans.com/907185247.
  2. To join via Polycom room system go to the IP Address: (bjn.vc) and enter the meeting ID: 907185247.
  3. To join via phone, use one of the following numbers and the Conference ID: 907185247.
    • US or Canada: +1 408 740 7256 or
    • US or Canada: +1 888 240 2560
  4. Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
  5. More information on connecting to bluejeans is available.

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Video-conference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947


Action Items

  1. None

Tentative Agenda

  1. Announcements
  2. Action Items
  3. Updates
  4. Any other business


Attendees: Elton, Mark D., Will M., Laura, Shayne (JLab), Andrei, Zisis (Regina)

  1. Announcements: none
  2. Action Items: We reviewed the list of student tasks drafted by Elton and Zisis (April 30 email).
    1. LEDs: they are operational. Need to operate them long term to check routine running and monitor gains; stability needs to be checked versus time. Some LEDs are not bright enough and need to run at a bias of 7V (instead of 6 or 6.4V). We should plan to run the LEDs at several biases and study these. Examining the LED amplitude versus time has not been automated yet, but we are getting close. Timing configuration adjustment is done manually. Automatic file sets are possible, ie choose which trigger to use without changing the cabling.
    2. Cosmics: need to start accumulating lots of stats by running nights and weekends. We have crude pedestals, that show a 2-3 channel drift. Pedestal can be looked at event-by-event. There are DAQ baseline channel drifts that occur in groups of channels (wedge-module or crate the culprit?). Needs further investigation. Andrei reported on the initial simulations giving Edep and Eattenuated. Will rerun. We will look at peak and mean values. Laura and Shayne have gotten the spectra from the bottom trigger scintillator (see Logbook Entry from June 5). We should try to take cosmics data over weekends. Runs so far have 20k events but we need much more. ROOT canvases need to be developed for each BCAL module (up and downstream). Elton will arrange for merging these into a single file, so we have 1 side of 1 module per page of plots. Note, upstream has higher event count rate than downstream -- top trigger counter is not located symmetrically, it is pushed more to the North side (N peak has 20 events, S is 8-10). Students will use xfig or equivalent to make drawing. Modules 45-48 down had a funny behavior that has been fixed: macro was not resetting all histos, all modules affected but 45 was worst. Modules 9-11 had high gain; Chris fixed the cabling causing this.
  3. Updates:
    1. System tests: 18 C runs are stable to +/- 0.1 C. New runs will be taken at 24 C and 12 C.
    2. Orlando/Yi are working on collecting all SiPM tests in the db. Need to use these to produce a comprehensive report and for inclusion into the TCR. Orlando leaves JLab to return to Chile on July 6. Elton/Mark/Yi will sit down with him to do the handover of all info.
    3. Orlando's EPICS GUI is done.
    4. Next steps:
      1. Activate TDCs and debug cabling etc, a la FADCs. Will will make a similar status panel for checking TDCs as he did for FADCs.
      2. Will will examine pulse integral, pulse time etc. and look at Michel electrons towards calibration.
      3. Mark/David/Zisis/Andrei will look at ccdb work (we need to iron out the context).
  4. Any other business
    1. Longer term we need to plan calibration efforts using simulations, for pi0 and other physics particles. Pursue this in August.
    2. All documentation required for the reviews is complete (thanks Elton).